OpenStudy (donnie1999):

---MEDAL--- Suppose f(x)=x^3 Find the graph of f(x)-3

OpenStudy (usukidoll):

there's a graph transformation ... so if we don't have a parenthesis then we are either going up which is the + sign or going down which is the - sign

OpenStudy (donnie1999):

@LynFran

OpenStudy (usukidoll):

f(x) = x^3 is the original graph and we're going 3 units ... up or down ?

OpenStudy (donnie1999):

up

OpenStudy (usukidoll):

no up means that we have a + sign and we don't have that

OpenStudy (usukidoll):

original graph f(x) =x^3 we have a graph transformation of f(x)-3 ... we have a - sign, so we're not shifting up.... we're shifting ?????

OpenStudy (donnie1999):

???

OpenStudy (usukidoll):

what's the opposite of going up?

OpenStudy (donnie1999):

lol down

OpenStudy (usukidoll):

yesssssssssssssssss >:D so we're going down 3 units

OpenStudy (usukidoll):

are you required to graph the original and the graph transformation? or is that all?

OpenStudy (donnie1999):

i have no idea srry im slow

OpenStudy (usukidoll):

ok, so was your question just find the graph of f(x) -3 only ?

OpenStudy (usukidoll):

then... we're done since the question had f(x) = x^3 which is the original graph and the graph transformation is 3 units down.